# Energy-Efficient Designs for SIM-Based Broadcast MIMO Systems

### v1

@@ VI Convergence
Let us now prove the convergence of Algorithm 2. First, for a given $\boldsymbol{\phi}$, Algorithm 2 achieves monotonic convergence, which can be shown as follows. From (23), it holds that
@@
where (a) is a due to (23), (b) is due to the fact that $\mathbf{U}^{(n+1)}$ is the optimal solution to (28) and that the optimal objective is no less than the objective of a feasible point, and (c) is true because it is easy to check that $\bar{h}(\mathbf{U}^{(n)};\mathbf{U}^{(n)})=h(\mathbf{U}^{(n)})$, i.e. the equality in (23) occurs when $\mathbf{U}=\mathbf{U}^{(n)}$. Regarding (b) above, note again that since (28) is a concave-convex fractional program, Dinkelbach’s method is guaranteed to converge to an optimal solution to (28). Consequently, the sequence $\{g(\mathbf{U}^{(n)})\}$ increases monotonically to an optimal solution to (18), and thus, Algorithm 2 is able to compute an optimal solution to (14). Next, for given covariance matrices, the SIM phase shift is optimized by a standard projected gradient method, for which the convergence is guaranteed. Also, the projected gradient method always yields an improved solution. In other words, Algorithm 2 generates a non-decreasing objective sequence. Since the feasible sets for the convariance matrices and phase shifts are continuous, the objective sequence produced by Algorithm 2 is guaranteed to converge.
@@ VII Simulation Results
In this section, we evaluate the EE of the considered systems using proposed algorithms by means of Monte Carlo simulations. First, we compare the EE of the proposed algorithms and three benchmark schemes. The first benchmark scheme, referred to as LIN w/o SIM, is based on linear precoding without the presence of a SIM. In that case, the total power consumption is $P_{r}+N_{t}P_{c}+P_{0}+LNP_{s}.$ The second benchmark scheme, referred to as LIN w/o prec., does not employ digital precoding; instead data streams are fed directly to transmit antennas, while the phase shifts of the SIM meta-elements are optimized as described in Algorithm 4. The achievable rate for a single user in this scheme can be calculated using (9) and the identity $\mathbf{P}_{k}\mathbf{P}_{k}^{H}=(P_{\mathrm{max}}/KN_{t})\mathbf{I}$, which ensures that the power constraint (10) is satisfied. The last benchmark scheme, termed LIN w/o prec. red. RF, also does not include a SIM, but differs from the previous one by utilizing a reduced number of transmit antennas and, consequently, a reduced number of RF chains. More precisely, this scheme uses $KN_{r}$ active transmit antennas (i.e., RF chains), each transmitting an independent data stream. The remaining transmit antennas (i.e., RF chains) are inactive, resulting in a total power consumption of $P_{r}+KN_{r}P_{c}+P_{0}+LNP_{s}.$
@@
The channel matrix between the BS and user $k$ is modeled according to the spatially-correlated channel model as $\mathbf{G}_{k}=\bar{\mathbf{G}}_{k}\mathbf{R}_{\text{T}}^{1/2}\in\mathbb{C}^{N_{r}\times L}$, where $\bar{\mathbf{G}}_{k}\in\mathbb{C}^{N_{r}\times L}$ denotes the channel between the last SIM layer and the receiver, and follows a complex Gaussian distribution $\mathcal{CN}(0,\beta\mathbf{I})$. The free space path loss between the transmitter and the receiver, $\beta$, is given by $\beta(d)=\beta(d_{0})+10b\log_{10}(d/d_{0})$, where $\beta(d_{0})=20\log_{10}(4\pi d_{0}/\lambda)$ is the free space path loss at the reference distance $d_{0}$, $b$ is the path loss exponent, and $d$ is the distance between the BS and user $k$. Moreover, $\mathbf{R}_{\text{T}}\in\mathbb{C}^{L\times L}$ is the spatial correlation matrix of the SIM, with its elements defined according to [18, Eq. (14), (15)].
@@
In the following simulation setup, the parameters are set as follows: $\lambda=5\,\text{cm}$ (i.e., $f=6\,\text{GHz}$), $N_{t}=16$, $N_{r}=2$, $W=100\thinspace\text{kHz}$, $\beta=3.5$, $d_{0}=1\,\text{m}$, $L=4$, $K=4$ and $\sigma^{2}=-110\,\text{dB}$. Unless otherwise specified, the number of meta-elements per SIM layer, $N$, is 100. The BS antennas are placed in a planner array parallel to the $xy$-plane and the position of its midpoint is $(30\,\mathrm{m},0,0)$. The inter antenna separation of the BS antennas is $\lambda/2$ in both dimensions. The BS SIM layers are also placed parallel to the $xy$-plane, with the midpoint of the $l$-th layer located at $(30\,\mathrm{m},0,l\lambda/2)$. Moreover, the meta-elements in each SIM layer are uniformly placed in a square formation, where each meta-element has dimensions $\lambda/2\times\lambda/2$. It is assumed that all users’ ULAs are parallel to the $x$-axis and the midpoint of the $k$-th user’s ULA is positioned at $(x_{k},y_{k},z_{k})$. The users’ coordinates are randomly selected such that $x_{k}$ is drawn from a uniform distribution between 1.6 m and 2 m with a resolution of 1 cm, $y_{k}$ is drawn from a uniform distribution between $-20$ m and 20 m with a resolution of 0.5 m, and $z_{k}$ is drawn from a uniform distribution between 80 m and 120 m with a resolution of 0.5 m. The circuit power per RF chain is $P_{c}=30\text{\thinspace dBm}$ and the basic power consumption at the BS is $P_{0}=40\thinspace\text{dBm}$ [7, 27]. The power consumption of each SIM meta-element is $P_{s}=10\thinspace\text{dBm}$ [28, 29]. For the DPC method, we use $\epsilon_{1,D}=\epsilon_{2,D}=10^{-6}$ and for linear precoding $\epsilon_{1,L}=\epsilon_{2,L}=10^{-6}$. Regarding the gradient-based optimization methods, the initial step size value is 1000, $\rho=1/2$ and $\delta=10^{-3}$. All results are averaged over 200 independent channel realizations.
Figure 2: EE versus number meta-elements per SIM layer.
The convergence of the proposed algorithms for different number of SIM layers is shown in Fig. 1. In general, we can see that the number of iterations required for the algorithms to converge increases with the number of SIM layers (i.e., meta-elements). A similar, but much more pronounced, effect was previously observed in [20], where the authors optimized the SIM phase shifts on a layer-by-layer basis. Moreover, the EE does not change monotonically with $L$. For DPC, the EE at convergence is almost the same for $L=4$ and $L=8$, which is visibly larger in the case of linear precoding. This can be attributed to the saturation of the achievable sum-rate as the number of SIM layers increases [19, Fig. 4], coupled with the fact that power consumption scales linearly with the number of meta-elements.
In Fig. 2, we present the EE evaluated for different number meta-elements per SIM layer. Specifically, the numbers of meta-elements considered per SIM layer are 25, 49, 100 and 196. The EE of both the proposed schemes and the benchmark schemes without precoding (i.e., LIN w/o prec. and LIN w/o prec. red. RF) increases with the number of meta-elements per SIM layer. The increase rates of the corresponding EE gradually reduces with $N$. The reason is that for these schemes, the transmitter operates full power, and thus the EE increases in line with the achieved sum-rate, which follows a logarithm function. Among the benchmark schemes without precoding, the LIN w/o prec. red. RF scheme achieves significantly higher EE. This is partly due to lower power consumption, as some of its RF chains are inactive. Additionally, in the LIN w/o prec. scheme, each antenna simultaneously transmits data for all users, while in the LIN w/o prec. red. RF scheme, each active antenna transmits an independent data stream, which better suppresses the multi-user interference. Since the LIN w/o SIM scheme does not incorporate a SIM, its EE is independent of the number of SIM meta-elements. For a small number of meta-elements per SIM layer, this benchmark scheme has the largest EE, while for a higher number of meta-elements the EE of other schemes becomes larger. Finally, we observe that the scheme with DPC achieves a slightly higher EE than the linear precoding scheme, and that this EE difference increases with the number of SIM meta-elements.
@@
Figure 3: EE (blue solid lines) and achievable sum-rate (red dashed lines) versus number SIM layer for the constant number of 400 meta-elements.
In Fig. 3, we present the EE and the achievable sum-rate of the considered system versus the number of SIM layers, while maintaining a constant total of $400$ meta-elements. As observed, the EE and the achievable sum-rate of the proposed schemes do not change monotonically with the number of SIM layers, $L$. Both the EE and the achievable sum-rate increase when $L$ changes from 1 to 2, which is likely due to the enhanced beamforming capabilities offered by multi-layer structures. However, with $L$ increases further, both of performance metrics significantly decrease, potentially reaching levels comparable to those of the LIN w/o SIM scheme, a benchmark scheme that does not utilize a SIM. This effect can be explained by the following reasoning:as the number of meta-elements per SIM layer decreases, the beamforming gain of each individual layer also reduces. Additionally, the signal propagation between adjacent SIM layers can be viewed as a form of path loss, which increases with the number of SIM layers. These two facts contribute to the observed reduction in both EE and achievable sum-rate when the number of SIM layers is greater than two. A similar trend is observed in the benchmark schemes without precoding (i.e., LIN w/o prec. and LIN w/o prec. red. RF), where the EE and the achievable sum-rate also decrease as $L$ increases. Among these two scheme, the LIN w/o prec. red. RF scheme achieves much better system performance.
@@
The EE of the considered schemes for different number of users, $K$, is shown in Fig. 4. For a small $K$, the LIN w/o prec. red. RF scheme provides the best EE, primarily due to the low number of RF chains used in this scheme. However, as $K$ increases, the number of RF chains used by this scheme becomes comparable to the total number of RF chains in other schemes. In addition, the lack of capability of adjusting the amplitude of the transmitted signal prevents this scheme from effectively suppressing the multi-user interference [25]. As a result, the EE of this scheme reduces as $K$ increases. For the same reason, the EE of the LIN w/o prec. scheme also decreases with an increasing number of users. On the other hand, the EE of the proposed schemes and the LIN w/o SIM scheme increase with $K$, because of the presence of digital precoders that can reduce or even eliminate the multi-user interference, allowing them to exploit the multiuser diversity. Comparing the EE of the proposed schemes with that of the benchmark schemes, we can see that using a SIM in combination with digital precoding almost always provides the best EE, except in cases where $K$ is very small.
Figure 5: EE versus the maximum transmit power.
Next, we study how the EE varies with the maximum transmit power, as shown in Fig. 5. The EE curves for all schemes exhibit an approximately logarithmic shape due to the logarithmic increase in the achievable rate. As expected, the proposed schemes achieve higher EE compared to benchmark schemes. Among all the benchmark schemes, the LIN w/o prec. red. RF scheme obtains the largest EE because of a smaller number of RF chains used. The difference between the EE of the proposed scheme with linear precoding scheme and that of any other benchmark scheme generally increases with transmit power, although these differences are almost stabilized at higher transmit power levels. Additionally, the DPC-based scheme consistently shows a noticeable improvement in EE over all other schemes, which can be attributed to the superior interference suppression capabilities of DPC.
In order to better understand the impact of realistic imperfections, we present the EE of the proposed schemes for the case of discrete SIM phase shifts in Fig. 6. The EE generally deteriorates as the number of quantization bits decreases. This effect is more evident for SIMs with a larger number of layers, since they contain more meta-elements and thus can cause a larger EE reduction. As a rule of thumb, at least 3 bits per meta-element are required for SIMs with a small $L$ (e.g., 2 or 4) to ensure that the EE reduction caused by quantization errors remains within acceptable limits. For SIMs with a larger $L$, the minimum number of bits per meta-element is expected to be higher.
@@
The per-iteration computational complexity of the proposed optimization schemes are shown in Table I. The relevant iteration counts for Dinkelbach’s method $I_{U}$ and $I_{X}$, and the number of the line search steps, $I_{\phi,D}$ and $I_{\phi,L}$, are averaged over the number of iterations required for each optimization scheme to reach 95 % of the EE at the 1000-th iteration. It can be observed that the number of iterations of Dinkelbach’s method remains almost unchanged as $K$ varies for both schemes, and the same holds true for the number of line search steps. Moreover, the computational complexity of the proposed optimization scheme for DPC-based systems is higher than that of the scheme with linear precoding when $K$ is 4 and 8. However, when $K$ increases to 12, the DPC-based scheme exhibits significantly lower complexity compared to the linear precoding scheme. This substantial increase in the complexity of the linear precoding scheme is due to the fact that the complexity of the precoding matrix optimization, $C_{x}$, is proportional to $N_{t}$ when $K=12$.
DPC Linear precoding
$K$ $I_{U}$ $I_{\phi,D}$ $C_{\text{DPC}}$ $I_{X}$ $I_{\phi,L}$ $C_{\text{LIN}}$
4 55 1 4367616 51 1 4129024
8 55 1 4480256 50 1 4409600
12 55 1 4592896 51 1 9947392
TABLE I: Comparison of the per-iteration computational complexities of the schemes with DPC and linear precoding.
VIII Conclusion

### v2

@@ VI Convergence
We now prove the convergence of Algorithm 2. First, for a given $\boldsymbol{\phi}$, Algorithm 2 achieves monotonic convergence, which can be shown as follows. From (20), it holds that
@@
where (a) follows from (20), (b) holds because $\mathbf{U}^{(n+1)}$ is the optimal solution to (25) and that the optimal objective is no less than the objective of a feasible point, and (c) is true because it is easy to check that $\bar{h}(\mathbf{U}^{(n)};\mathbf{U}^{(n)})=h(\mathbf{U}^{(n)})$, i.e. the equality in (20) occurs when $\mathbf{U}=\mathbf{U}^{(n)}$. Regarding (b) above, note again that since (25) is a concave-convex fractional program, Dinkelbach’s method is guaranteed to converge to its optimal solution. Consequently, the sequence $\{g(\mathbf{U}^{(n)})\}$ increases monotonically to an optimal solution to (17), and thus, Algorithm 2 is able to compute an optimal solution to (13). Next, for given covariance matrices, the SIM phase shift is optimized by a standard projected gradient method, for which the convergence is guaranteed. Also, the projected gradient method always yields an improved solution. In other words, Algorithm 2 generates a non-decreasing objective sequence. Since the feasible sets for the covariance matrices and phase shifts are continuous, the objective sequence produced by Algorithm 2 is guaranteed to converge.
@@ VII Simulation Results
In this section, we evaluate the EE of the proposed schemes with DPC and LP, which are denoted as SIM-DPC and SIM-LP, respectively, through Monte Carlo simulations and compare them with three benchmark schemes:
  1. 1.
    LP-NoSIM: The first benchmark scheme uses LP without SIM integration, with a total power consumption of $P_{t}+N_{t}P_{c}+P_{0}.$
  2. 2.
    SIM-NoLP: The second scheme, omits digital precoding, feeding data streams directly to transmit antennas while optimizing SIM phase shifts as in Algorithm 4. Let $N_{t,k}$ BS antennas are allocated for transmission to the $k$-th user. Since all $N_{t}$ transmit antennas are active, $N_{t,k}$ equals $\left\lfloor N_{t}/K\right\rfloor$ or $\left\lfloor N_{t}/K\right\rfloor+1$. Thus, each data stream of user $k$ is transmitted by $\left\lfloor N_{t,k}/N_{r}\right\rfloor$ or $\left\lfloor N_{t,k}/N_{r}\right\rfloor+1$ BS antennas.
  3. 3.
    SIM-NoLP-RedRF: The third scheme excludes LP but differs from SIM-NoLP by utilizing a reduced number of transmit antennas (i.e., RF chains). Specifically, it activates $KN_{r}$ transmit antennas (i.e., RF chains), each transmitting an independent data stream, while the remaining transmit antennas remain inactive. As such, the total power consumption of $P_{t}+KN_{r}P_{c}+P_{0}+LNP_{s}$. Note that for the case when $KN_{r}>N_{t}$, SIM-NoLP transmits signal to a subset of randomly chosen users, while LP-NoSIM-RedRF cannot operate. To ensure fair comparison, data for SIM-NoLP and LP-NoSIM-RedRF are scaled by $\sqrt{P_{\max}/N_{t}}$ and $\sqrt{P_{\max}/KN_{r}},$ respectively.
Channel Modeling
@@
The channel matrix between the BS and user $k$ is modeled as $\mathbf{G}_{k}=\bar{\mathbf{G}}_{k}\mathbf{R}_{\text{T}}^{1/2}\in\mathbb{C}^{N_{r}\times L}$, where $\bar{\mathbf{G}}_{k}\in\mathbb{C}^{N_{r}\times L}$ denotes the channel between the last SIM layer and the receiver, distributed as $\mathcal{CN}(0,\beta\mathbf{I})$. The path-loss between the BS and user $k$ is given by $\beta(d_{k})=\beta(d_{0})+10b\log_{10}(d_{k}/d_{0})$, where $\beta(d_{0})=20\log_{10}(4\pi d_{0}/\lambda)$ is the free space path-loss at the reference distance $d_{0}$, $b$ is the path-loss exponent, and $d_{k}$ is the distance between the BS and user $k$. Moreover, $\mathbf{R}_{\text{T}}\in\mathbb{C}^{L\times L}$ is the spatial correlation matrix of the SIM, defined as in [17, Eq. (14), (15)].
System Parameters
@@
In the simulation setup, parameters are set as follows: $\lambda=5\,\text{cm}$ (i.e., $f=6\,\text{GHz}$), $N_{t}=16$, $N_{r}=2$, $W=100\thinspace\text{kHz}$, $b=3.5$, $d_{0}=1\,\text{m}$, $L=4$, $K=4$, $P_{\max}=5\,\mathrm{W}$ and $\sigma^{2}=-110\,\text{dB}$. Unless specified otherwise, the number of meta-elements per SIM layer, $N$, is 100. The BS antennas are placed in a planar array parallel to the $xy$-plane, with the midpoint at $(30\,\mathrm{m},0,0)$ and inter-antenna spacing of $\lambda/2$. SIM layers are also placed parallel to the $xy$-plane, with the midpoint of the $l$-th layer at $(30\,\mathrm{m},0,l\lambda/2)$. Meta-elements, each with dimension $\lambda/2\times\lambda/2$, are uniformly placed in a square grid on each SIM layer. Hence, any mutual coupling between different meta-elements can be neglected. Users’ ULAs are parallel to the $x$-axis, with the midpoint of the $k$-th user’s ULA at $(x_{k},y_{k},z_{k})$. Users’ coordinates are uniformly distributed such that $x_{k}\in[1.6\,\text{m},2\,\text{m}],y_{k}\in[-20\,\text{m},20\,\text{m}],z_{k}\in[80\,\text{m},120\,\text{m}]$. The circuit power per RF chain is $P_{c}=30\text{\thinspace dBm},$ the basic power consumption at the BS is $P_{0}=40\thinspace\text{dBm}$ [7, 26], and the power consumption per SIM meta-element is $P_{s}=10\thinspace\text{dBm}$ [27, 28]. Optimization parameters are set $\epsilon_{D}=10^{-6}$ for DPC, and $\epsilon_{1,L}=\epsilon_{2,L}=10^{-6}$ for LP. Gradient-based methods use an initial step size of 1000, $\rho=1/2$ and $\delta=10^{-3}$. Results are averaged over 200 independent channel realizations.
Initialization
For DPC the initial $\mathbf{S}_{k}$ matrices are randomly generated Hermitian matrices satisfying (13b) with equality. For LP the initial $\mathbf{X}_{k}$ are randomly generated matrices satisfying (42b) with equality. The initial elements of $\mathbf{\boldsymbol{\phi}}$ are randomly chosen from the unit circle.
Figure 3: $E$ versus number meta-elements per SIM layer.
The convergence of the proposed algorithms for different numbers of SIM layers is shown in Fig. 2. In general, we can see that the number of iterations required to converge increases with the number of SIM layers (i.e., meta-elements), consistent with previous observations in [18], where a layer-by-layer approach for SIM optimization was used. Notably, the gap in EE performance between SIM-DPC and SIM-LP remains small for any number of SIM layers, making SIM-LP a practical and efficient solution for SIM-based systems. Moreover, the EE is not monotonic with $L$. For SIM-DPC, the EE at convergence is almost the same for $L=4$ and $L=8$, while for SIM-LP, it is slightly higher when $L=8$. This behavior results from the fact that the total power consumption scales linearly with the number of meta-elements, while the achievable sum-rate tends to be saturated [29, Fig. 4].
Fig. 3 shows the EE for different number meta-elements per SIM layer ($N=25,49,100,196)$. The EE of both proposed and the non-precoding benchmark schemes (i.e., SIM-NoLP and SIM-NoLP-RedRF) increases with $N$ but the increase rate gradually declines due to the logarithmic behavior of the achievable sum-rate at near-full transmit power. Among the non-precoding benchmarks, SIM-NoLP-RedRF achieves higher EE than SIM-NoLP due to reduced power consumption from inactive RF chains. Moreover, SIM-NoLP suffers from increased multi-stream interference as in this case each data stream is transmitted from two different transmit antennas without digital precoding, which potentially worsens as $N$ increases, further enlarging the EE gap between these benchmarks. On the other hand, the saturation of the achievable sum-rate, and consequently the EE, with the further increase of the number of meta-elements results in almost the same enlargement of the EE for the two benchmark schemes. For a small $N$, LP-NoSIM achieves the largest EE, but is eventually inferior as other schemes benefit from an increased number of meta-elements. Finally, SIM-DPC achieves slightly higher EE than SIM-LP, with the gap increasing as $N$ grows.
@@
Figure 4: $E$ (blue solid lines) and achievable sum-rate (red dashed lines) versus number SIM layer for the constant number of 400 meta-elements.
Fig. 4 plots the EE and the achievable sum-rate of the considered system versus the number of SIM layers, while keeping a total of $400$ meta-elements. Importantly, we find that SIM-LP is as good as SIM-DPC for both performance metrics, regardless of the distribution of meta-elements on SIM layers. and neither metric is monotonic with $L$. Specifically, both increase from $L=1$ to $L=2$, which is apparently due to the enhanced beamforming capabilities offered by multi-layer structures. However, with $L$ increases further, both performance metrics significantly decrease, potentially approaching LP-NoSIM. This occurs because fewer meta-elements per layer reduce the beamforming capability of each layer, while inter-layer propagation introduces additional path-loss as the number of SIM layers increases. A similar trend is observed in non-precoding benchmarks (i.e., SIM-NoLP and SIM-NoLP-RedRF), where the EE and the achievable sum-rate also decrease for $L>4$. Among them, SIM-NoLP-RedRF achieves higher EE, despite a lower achievable sum-rate as it operates with fewer RF chains.
@@
Fig. 5 demonstrates the EE for different number of users, $K$. For a small $K$, SIM-NoLP-RedRF provides the highest EE due to its reduced RF chain (and thus power) usage. However, as $K$ increases, its EE declines since the number of RF chains becomes comparable to other schemes, and the lack of amplitude control of the transmitted signal limits its multi-user interference suppression capability [23]. On the other hand, the EE of SIM-NoLP shows a non-monotonic behavior with $K$. Specifically, its EE slightly improves from $K=2$ to $K=4$ due to the SIM being able to exploit multi-user diversity gains. However, when $K$ further increases, the EE of both SIM-NoLP and SIM-NoLP-RedRF declines. The EE of the proposed SIM-based schemes and LP-NoSIM increase with $K$ since digital precoding can effectively mitigate the multi-user interference, allowing them to exploit the multi-user diversity. Moreover, SIM-DPC achieves slightly higher EE than SIM-LP, and the gap widens as $K$ increases. This observation further supports LP as an efficient, practical solution for SIM-based systems. Overall, we conclude that integrating SIM with digital precoding consistently yield the best EE, except for very small $K$.
Figure 6: $E$ versus the maximum transmit power.
Next, Fig. 6 illustrates how the EE varies with the maximum transmit power. All schemes in consideration exhibit a logarithmic EE increase due to the logarithmic growth of the achievable rate. At low transmit power, the proposed schemes (i.e. SIM-DPC and SIM-LP) achieve the same EE, as do the benchmark schemes. As the transmit power increases, SIM-DPC consistently shows a more noticeable improvement in EE over all other schemes, which is attributed to the superior interference suppression capabilities of DPC. As expected, the proposed schemes outperform benchmark schemes, with, SIM-NoNP-RedRF obtaining the highest EE among them because of fewer RF chains being used. The difference in the EE between SIM-LP and other benchmark schemes increases with transmit power stabilizes at higher transmit power levels.
To assess the impact of realistic imperfections, Fig. 7 presents the EE of the proposed schemes for discrete SIM phase shifts. The EE generally deteriorates as the number of quantization bits decreases, especially for SIMs with a larger number of layers since the quantization errors are amplified. As a rule of thumb, at least 3 bits per meta-element are required for SIMs for a small $L$ (e.g., 2 or 4) to maintain acceptable EE levels, while larger $L$ requires finer resolution.
@@
Finally, Fig. 8 shows the per-iteration computational complexity of the proposed optimization schemes versus the number of transmit antennas. The LP-NoSIM scheme achieves the lowest complexity compared to other schemes, as precoding without SIM involves much fewer elements than SIM meta-elements. Due to the same reason, the complexity of the benchmark schemes without precoding is very similar to that of the proposed schemes, and it is not even possible to observe any deference between DPC and LP.
Figure 8: Computational complexity versus the number transmit antennas.
Figure 9: Computational complexities of the proposed schemes with DPC and LP that are without the SIM, and the scheme from [19] versus the number transmit antennas.
To better highlight the difference between DPC and LP, we compare the computational complexities of their appropriate schemes without SIM (i.e., DPC-NoSIM and LP-NoSIM) in Fig. 9. As a benchmark, we consider the DPC-based scheme from [19], also without SIM, referred to as DPC-bench. Its complexity is $O(I_{D}K(N_{t}^{2}N_{r}+N_{t}^{3}))$, where $I_{D}$ is the required number of loops. It can be observed that LP-NoSIM has the highest complexity for $N_{t}=6$ and $N_{t}<KN_{r}$. As $N_{t}$ further increases, the complexities of DPC-NoSIM and DPC-bench become larger than that of LP-NoSIM. This is because the LP-NoSIM complexity remains approximately constant for $N_{t}\geq KN_{r}$, as indicated by (57)). In addition, DPC-bench has a higher complexity than DPC-NoSIM, despite both achieving the same energy efficiency.
VIII Conclusion, Insights, and Future Work